ou need to accumulate $10,000. To do so, you plan to make deposits of $1,150 per year - with the first payment being made a year from today - into a bank account that pays 8.62% annual interest. Your last deposit will be less than $1,150 if less is needed to round out to $10,000. How many years will it take you to reach your $10,000 goal? Round your answer up to the nearest whole number. 7.year(s)

How large will the last deposit be? Round your answer to the nearest cent.
Answer #1

FV = PMT*((1+r)n-1)/r
10,000 = 1150 *((1+8.62%)n-1)/8.62%
1.0862n =(1 + 10000*8.62%/1150)
Applying log on both sides = log 1.0862n = log 1.74956
n = log 1.74596 /log 1.0862= 6.77 years

FV of annuity in 6 years = 1150*((1+8.62%6-1)/8.62% = 8569.29
Value of 10,000 in year 6 = 10,000/1.0862 = 9206.47

The last deposit value = 9206.47 - 8569.29 = 637.12

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well